# Detection of Gravitational Anomaly at Low Acceleration from a Highest-quality Sample of 36 Wide Binaries with Accurate 3D Velocities

### v1

@@ E.5 Speckle Holography
Gaia DR3 2021978621937192832 2025.4371 1.04 0.60
Gaia DR3 2021979584009866368 2025.4371 0.42 0.28
Gaia DR3 2021979588354850176 2025.4371 2.05 1.73
Gaia DR3 2029432043779954432 2025.4400 0.83 0.45
Gaia DR3 2029433521248546304 2025.4400 0.10 0.09
Gaia DR3 2032121960300897280 2025.4398 1.10 0.61
Gaia DR3 2047012822409354752 2025.4399 0.52 0.32
Gaia DR3 2047014368597597696 2025.4399 0.45 0.29
Gaia DR3 2051915235490493952 2025.4399 0.30 0.22
Gaia DR3 2051915304209972992 2025.4399 0.30 0.21
Gaia DR3 2056444914159643008 2025.4426 0.49 0.31
Gaia DR3 2056444914159643136 2024.6781 0.50 0.33
Gaia DR3 2056444948519384064 2025.4426 0.51 0.33
Gaia DR3 2062485390533084544 2024.6780 0.30 0.20
Gaia DR3 2062485390533085568 2024.6780 0.32 0.23
Gaia DR3 2066763860841212672 2024.6783 1.48 1.12
Gaia DR3 2066763865140783104 2024.6783 2.93 2.51
Gaia DR3 2067274450853577344 2025.4455 0.70 0.38
Gaia DR3 2067274515272929536 2025.4427 0.29 0.20
Gaia DR3 2076871091425583232 2025.4399 1.29 0.88
Gaia DR3 2076871091425586944 2025.4399 0.91 0.51
Gaia DR3 2078886285778453632 2025.4400 0.73 0.41
Gaia DR3 2078886290079138560 2025.4400 0.59 0.36
Gaia DR3 2079604404307344128 2025.4372 0.20 0.14
Gaia DR3 2079604442969955968 2025.4372 0.82 0.49
Gaia DR3 2086713404118555520 2025.4426 0.37 0.26
Gaia DR3 2086713404118555648 2025.4426 0.28 0.20
Gaia DR3 2087586175833524480 2024.6780 0.35 0.25
Gaia DR3 2087586175833525376 2024.6780 0.23 0.14
Gaia DR3 2091495764301335424 2025.4452 0.11 0.10
Gaia DR3 2091495798661081344 2025.4452 0.13 0.11
Gaia DR3 2098996731404747776 2025.4399 0.53 0.33
Gaia DR3 2098996731404747904 2025.4399 1.47 1.02
Gaia DR3 2098996804424367360 2025.4399 0.37 0.25
Gaia DR3 2099859268216511488 2025.4399 1.14 0.77
Gaia DR3 2099859268216512000 2025.4399 1.41 0.90
Gaia DR3 2100011516217168640 2025.4399 1.46 1.03
Gaia DR3 2100011516217169920 2025.4399 1.49 1.00
Gaia DR3 2100451630105040256 2025.4399 1.26 0.83
Gaia DR3 2100451630105041152 2025.4399 0.38 0.26
Gaia DR3 2100501760964021504 2025.4399 0.37 0.25
Gaia DR3 2100502001482188032 2025.4399 0.41 0.28
Gaia DR3 2104998900960077056 2025.4398 0.00 0.00
Gaia DR3 2104999343337767424 2025.4398 0.60 0.35
Gaia DR3 211413206732904960 2024.8650 0.34 0.23
Gaia DR3 211413206732905600 2024.8650 0.37 0.26
Gaia DR3 2116748243555689728 2025.4397 0.48 0.32
Gaia DR3 2116748247853607168 2025.4397 0.26 0.18
Gaia DR3 2116892318236994816 2025.4398 0.22 0.16
This is a triple system. Gaia DR3 2119132916774654080 2025.4398 0.36 0.25
@@
This is a triple system. Gaia DR3 2119132916774654208 2025.4398 1.11 0.73
@@
Gaia DR3 2126777988630768000 2025.4400 0.75 0.44
Gaia DR3 2126777992924832128 2025.4400 1.00 0.65
Gaia DR3 2128304179486008704 2025.4426 0.37 0.26
Gaia DR3 2128304183783972992 2025.4426 0.28 0.20
Gaia DR3 2129540099271049216 2025.4426 0.82 0.46
Gaia DR3 2129540103571086592 2025.4426 0.99 0.56
Gaia DR3 2130903047014230528 2025.4426 0.63 0.37
Gaia DR3 2130903223113004032 2025.4426 0.33 0.23
Gaia DR3 2137630172689997824 2025.4372 0.31 0.21
Gaia DR3 2137630172690002176 2025.4372 0.29 0.20
Gaia DR3 2140767319879292160 2025.4372 0.39 0.28
Gaia DR3 2140767560397457536 2025.4453 0.29 0.21
Gaia DR3 2148617626742496512 2025.4371 0.29 0.20
Gaia DR3 2148617729821714688 2025.4371 0.43 0.29
Gaia DR3 2157622725756418816 2025.4398 0.29 0.21
Gaia DR3 2157623451608162688 2025.4398 0.88 0.47
Gaia DR3 2171870747199482496 2024.6864 1.03 0.59
Gaia DR3 2171870751495498752 2024.6864 0.67 0.38
Gaia DR3 2179261496856372224 2024.6864 0.93 0.58
Gaia DR3 2179261496856372352 2024.6864 0.92 0.56
Gaia DR3 2182397334073261568 2024.6782 1.03 0.62
Gaia DR3 2182397338377483264 2024.6782 0.84 0.47
Gaia DR3 2186439383632375296 2024.6863 2.40 1.77
Gaia DR3 2186439482412215808 2024.6863 0.62 0.40
Gaia DR3 219593745044391552 2024.8618 0.37 0.26
Gaia DR3 219605599154126976 2024.8618 0.49 0.32
Gaia DR3 2196841961921480960 2024.6864 0.86 0.53
Gaia DR3 2196841966222594944 2024.6864 1.75 1.40
Gaia DR3 2201661091331626752 2024.6867 0.25 0.18
Gaia DR3 2201661297490052096 2024.6867 0.24 0.17
Gaia DR3 2201834466266276864 2024.6867 0.23 0.16
Gaia DR3 2201834470572701952 2024.6867 0.28 0.20
Gaia DR3 222299432702445184 2024.8618 1.45 1.07
Gaia DR3 222299436999652096 2024.8618 1.48 1.13
Gaia DR3 2242465891977167872 2025.4426 0.24 0.17
Gaia DR3 2242465891977204992 2025.4426 0.43 0.30
Gaia DR3 229737942400057728 2024.8618 0.59 0.36
Gaia DR3 229737942400695424 2024.8618 1.75 1.35
Gaia DR3 2503402743097074304 2024.8618 1.01 0.58
Gaia DR3 2503402747391938688 2024.8618 0.83 0.45
Gaia DR3 2512615551025883136 2024.6868 0.95 0.54
Gaia DR3 2512615555321101696 2024.6868 0.31 0.22
Gaia DR3 2514529598906714880 2024.8618 0.52 0.33
Too weak Gaia DR3 2514541414361083648 2024.8618 >2.>2.. >2.>2.
Gaia DR3 2524864836409161728 2024.6868 0.36 0.26
Gaia DR3 2529474126591876096 2024.6815 0.41 0.28
Gaia DR3 2529661799483080192 2024.6815 0.29 0.21
Gaia DR3 2531509253895359744 2024.6868 0.23 0.17
Gaia DR3 2531509253895360000 2024.6868 0.29 0.21
Gaia DR3 2555905080453468544 2024.6815 0.30 0.21
Gaia DR3 2555905080453468800 2024.6815 0.26 0.19
Gaia DR3 2556421335522520576 2024.6816 0.29 0.21
Gaia DR3 2556421610401138176 2024.6816 0.44 0.30
Gaia DR3 2565584802867037696 2024.6868 0.19 0.13
Gaia DR3 2565584837226776576 2024.6868 0.21 0.14
Gaia DR3 2572437745309855488 2024.6869 0.23 0.14
Gaia DR3 2572437749605558272 2024.6869 0.35 0.24
Gaia DR3 2573278051366910336 2024.6869 0.16 0.12
Gaia DR3 2573278120086386432 2024.6869 0.16 0.12
Triple system, variable Gaia DR3 2581429521337662464 2024.6868 0.26 0.18
Triple system, variable Gaia DR3 2581429521338139264 2024.6868 0.28 0.20
Gaia DR3 2586004584926025600 2024.6869 0.30 0.21
Gaia DR3 2586004589220994560 2024.6869 0.28 0.20
Gaia DR3 2611384547405130752 2024.6840 0.34 0.24
Gaia DR3 2611384547405183488 2024.6840 0.32 0.23
Gaia DR3 2621082995876646400 2024.6840 0.31 0.22
Gaia DR3 2621082995876646528 2024.6840 0.37 0.26
Gaia DR3 2638615189818036608 2024.6840 0.13 0.11
Gaia DR3 2638615189818036736 2024.6840 0.13 0.11
Gaia DR3 2642922251741817344 2024.6840 0.51 0.32
Gaia DR3 2642922286101533696 2024.6840 0.56 0.34
Gaia DR3 266993755737826176 2024.8650 1.07 0.65
Gaia DR3 266993755740665600 2024.8650 1.09 0.68
Gaia DR3 2683565561622947584 2024.6839 0.48 0.32
Gaia DR3 2683565630342923520 2024.6839 0.24 0.17
Gaia DR3 2688137537130343936 2024.6838 0.17 0.12
Gaia DR3 2688137537130451456 2024.6838 0.22 0.16
Gaia DR3 2722918766408972288 2024.6839 0.34 0.24
Gaia DR3 2722918796473758720 2024.6839 0.38 0.26
Gaia DR3 2725049173267026560 2024.6838 0.17 0.12
Gaia DR3 2725049173267027968 2024.6838 0.29 0.21
Gaia DR3 2731572610114153856 2024.6840 0.19 0.14
Gaia DR3 2731572610114154112 2024.6840 0.24 0.18
Gaia DR3 2738818666619511808 2024.6868 0.32 0.22
Gaia DR3 2738818666619512064 2024.6868 0.35 0.24
Gaia DR3 2739496687336986112 2024.6841 0.20 0.14
Gaia DR3 2739498199165473792 2024.6841 0.25 0.18
Gaia DR3 2741770924059034624 2024.6814 0.48 0.30
Gaia DR3 2741770958420224000 2024.6814 0.21 0.15
Gaia DR3 2743364563085132800 2024.6841 1.73 1.42
Gaia DR3 2743364563085132928 2024.6841 1.77 1.44
Gaia DR3 2752042939643757952 2024.6868 0.23 0.16
Gaia DR3 2752042939643758080 2024.6868 0.22 0.15
Gaia DR3 2763789469038658432 2024.6841 0.23 0.16
Gaia DR3 2763789469039961472 2024.6841 0.17 0.12
Gaia DR3 2799550942998190592 2024.6815 0.00 0.00
Gaia DR3 2799550947293531648 2024.6814 0.42 0.28
Gaia DR3 280403536989156224 2024.8650 0.32 0.23
Too weak Gaia DR3 280403541287570176 2024.8650 >2.>2. >2.>2.
Gaia DR3 2813290096701231104 2024.6840 0.77 0.39
Gaia DR3 2813290100998398464 2024.6840 0.75 0.40
Gaia DR3 2833895636096202752 2024.6840 0.32 0.22
Gaia DR3 2833895636096692224 2024.6757 0.28 0.21
Gaia DR3 2836313908842872704 2024.6866 0.29 0.19
Gaia DR3 2836313908842872832 2024.6866 0.25 0.17
Gaia DR3 2837991381336616320 2024.6867 0.45 0.31
Gaia DR3 2837991454349643136 2024.6867 0.45 0.30
Gaia DR3 2847357575430420992 2024.6815 0.26 0.18
Gaia DR3 2847357575430421120 2024.6815 0.65 0.38
Gaia DR3 2849031405789315328 2024.6814 0.20 0.13
Gaia DR3 2849031410085904640 2024.6814 0.33 0.23
Gaia DR3 2860343426229962752 2024.6814 0.24 0.16
Gaia DR3 2865611564396936192 2024.6866 0.29 0.21
Gaia DR3 2865611564396936320 2024.6866 0.35 0.24
Gaia DR3 2868784063464098048 2024.6866 0.78 0.49
Gaia DR3 2868784136478476032 2024.6866 0.20 0.13
Gaia DR3 2874196929143032192 2024.6815 1.22 0.83
Gaia DR3 2874196929143032448 2024.6815 1.20 0.77
Gaia DR3 2879160021551346176 2024.6866 0.52 0.34
Gaia DR3 2879160021551346688 2024.6866 0.43 0.29
Gaia DR3 293553975230799104 2024.6869 0.31 0.21
Gaia DR3 293554074014451200 2024.6869 0.21 0.13
Gaia DR3 2989629982019538304 2024.8621 2.52 1.85
Gaia DR3 2989630016379276160 2024.8621 2.23 1.86
Gaia DR3 3002343871195786496 2024.8622 0.38 0.27
Gaia DR3 3002531548383047296 2024.8622 0.37 0.26
Gaia DR3 3015129546453825920 2024.8622 1.38 0.95
Gaia DR3 3015130405447132288 2024.8621 1.34 0.89
Gaia DR3 3033070586928188416 2024.8624 0.47 0.32
Gaia DR3 3033070655647665536 2024.8624 1.47 1.16
Gaia DR3 3033419235197542912 2024.8624 0.74 0.43
Gaia DR3 3033442909056369792 2024.8624 0.63 0.39
Gaia DR3 3033693700785580800 2024.8624 0.95 0.56
@@
Gaia DR3 3035042522369773824 2024.8624 1.27 0.87
Gaia DR3 3035042526672692096 2024.8624 1.12 0.83
Gaia DR3 3042299715729341056 2024.8624 1.03 0.63
Gaia DR3 3042300093686461312 2024.8624 0.95 0.57
Gaia DR3 3046204150242468480 2025.2173 0.58 0.37
Gaia DR3 3046204180298475264 2025.2173 0.43 0.30
Gaia DR3 3048356684771809536 2024.8624 0.37 0.27
@@
Gaia DR3 3061488839331542912 2024.8624 0.90 0.52
Gaia DR3 3061488839334653696 2024.8624 0.91 0.54
Gaia DR3 3075884367113894400 2024.8625 1.57 1.30
Gaia DR3 3075884367113894528 2024.8625 1.22 0.83
Gaia DR3 3079739250589041024 2024.8625 1.70 1.30
Gaia DR3 3079739254881696256 2024.8625 1.18 0.68
Gaia DR3 3084652044636398976 2024.8625 0.35 0.25
Gaia DR3 3084652044636399104 2024.8625 0.38 0.27
Gaia DR3 3103862814877943296 2024.8622 0.38 0.27
Gaia DR3 3103862814877944064 2025.2174 0.45 0.31
Gaia DR3 3107726150841412224 2024.8624 1.54 1.31
Gaia DR3 3107726155141310080 2024.8624 1.52 1.25
Gaia DR3 3155908468358227968 2024.8624 1.20 0.85
Gaia DR3 3155908575735398400 2024.8624 1.01 0.67
Gaia DR3 3158878734598549376 2025.2173 0.48 0.31
Gaia DR3 3158878734598549504 2024.8623 0.63 0.37
Gaia DR3 3158926322836178816 2025.2173 0.42 0.28
Gaia DR3 3158926322836178944 2024.8623 0.95 0.49
Gaia DR3 3170300942420466176 2025.2173 0.28 0.20
Gaia DR3 3170303755624614528 2024.8623 0.48 0.31
Gaia DR3 3170303759920420224 2024.8623 0.33 0.23
Gaia DR3 3170394607068638336 2025.2173 0.30 0.22
Gaia DR3 3174712286151938944 2024.8621 1.19 0.80
Gaia DR3 3174712354871414912 2024.8621 0.97 0.58
Gaia DR3 3194331181363949312 2024.8619 0.96 0.55
Gaia DR3 3194331284443164288 2024.8619 0.86 0.48
@@
Gaia DR3 3194720889516362880 2024.8619 1.08 0.72
Gaia DR3 3201928153876092416 2024.8620 1.78 1.46
Gaia DR3 3201928153876683648 2024.8620 1.97 1.60
Gaia DR3 3208097204381858304 2024.8622 0.78 0.43
Gaia DR3 3208097582339739136 2024.8622 0.71 0.40
Gaia DR3 3208920321978859008 2024.8622 0.43 0.30
Gaia DR3 3208920326274769664 2024.8622 1.48 1.07
@@
Gaia DR3 3211638593896898944 2024.8622 0.42 0.28
Gaia DR3 3213811602532945152 2024.8622 1.48 1.13
Gaia DR3 3213811641189319680 2024.8622 1.52 1.14
Gaia DR3 3216741186843662720 2024.8622 2.00 1.60
Gaia DR3 3216741186843662848 2024.8622 2.88 2.44
@@
Gaia DR3 3223175082210856192 2024.8622 1.37 0.88
Gaia DR3 3233274852427241216 2024.8620 0.51 0.33
Too weak Gaia DR3 3233274955506455936 2024.8620 >2.>2. >2.>2.
Gaia DR3 3237157051889659904 2024.8622 0.70 0.40
Gaia DR3 3237157159265190400 2024.8622 0.83 0.46
Gaia DR3 3239798461074336512 2024.8649 0.68 0.39
Gaia DR3 3239798461074336640 2024.8649 0.79 0.43
Gaia DR3 3242368702645853184 2024.8649 0.83 0.45
Gaia DR3 3242368706941778688 2024.8649 0.77 0.43
Gaia DR3 3251884739561945472 2024.8620 2.90 2.34
Gaia DR3 3251884739561945600 2024.8620 0.93 0.54
Gaia DR3 3258912371210481152 2024.8620 0.72 0.41
Gaia DR3 3258912439929957376 2024.8620 1.19 0.83
Gaia DR3 3266980170921153920 2024.8618 1.18 0.74

### v2

@@ E.5 Speckle Holography
Gaia DR3 2021978621937192832 2025.4371 0.520 0.300
Gaia DR3 2021979584009866368 2025.4371 0.210 0.140
Gaia DR3 2021979588354850176 2025.4371 1.025 0.865
Gaia DR3 2029432043779954432 2025.44 0.415 0.225
Gaia DR3 2029433521248546304 2025.44 0.050 0.045
Gaia DR3 2032121960300897280 2025.4398 0.550 0.305
Gaia DR3 2047012822409354752 2025.4399 0.260 0.160
Gaia DR3 2047014368597597696 2025.4399 0.225 0.145
Gaia DR3 2051915235490493952 2025.4399 0.150 0.110
Gaia DR3 2051915304209972992 2025.4399 0.150 0.105
Gaia DR3 2056444914159643008 2025.4426 0.245 0.155
Gaia DR3 2056444914159643136 2024.6781 0.250 0.165
Gaia DR3 2056444948519384064 2025.4426 0.255 0.165
Gaia DR3 2062485390533084544 2024.678 0.150 0.100
Gaia DR3 2062485390533085568 2024.678 0.160 0.115
Gaia DR3 2066763860841212672 2024.6783 0.740 0.560
Gaia DR3 2066763865140783104 2024.6783 1.465 1.255
Gaia DR3 2067274450853577344 2025.4455 0.350 0.190
Gaia DR3 2067274515272929536 2025.4427 0.145 0.100
Gaia DR3 2076871091425583232 2025.4399 0.645 0.440
Gaia DR3 2076871091425586944 2025.4399 0.455 0.255
Gaia DR3 2078886285778453632 2025.44 0.365 0.205
Gaia DR3 2078886290079138560 2025.44 0.295 0.180
Gaia DR3 2079604404307344128 2025.4372 0.100 0.070
Gaia DR3 2079604442969955968 2025.4372 0.410 0.245
Gaia DR3 2086713404118555520 2025.4426 0.185 0.130
Gaia DR3 2086713404118555648 2025.4426 0.140 0.100
Gaia DR3 2087586175833524480 2024.678 0.175 0.125
Gaia DR3 2087586175833525376 2024.678 0.115 0.070
Gaia DR3 2091495764301335424 2025.4452 0.055 0.050
Gaia DR3 2091495798661081344 2025.4452 0.065 0.055
Gaia DR3 2098996731404747776 2025.4399 0.265 0.165
Gaia DR3 2098996731404747904 2025.4399 0.735 0.510
Gaia DR3 2098996804424367360 2025.4399 0.185 0.125
Gaia DR3 2099859268216511488 2025.4399 0.570 0.385
Gaia DR3 2099859268216512000 2025.4399 0.705 0.450
Gaia DR3 2100011516217168640 2025.4399 0.730 0.515
Gaia DR3 2100011516217169920 2025.4399 0.745 0.500
Gaia DR3 2100451630105040256 2025.4399 0.630 0.415
Gaia DR3 2100451630105041152 2025.4399 0.190 0.130
Gaia DR3 2100501760964021504 2025.4399 0.185 0.125
Gaia DR3 2100502001482188032 2025.4399 0.205 0.140
Too weak Gaia DR3 2104998900960077056 2025.4398 >2.>2. >2.>2.
Gaia DR3 2104999343337767424 2025.4398 0.300 0.175
Gaia DR3 211413206732904960 2024.865 0.170 0.115
Gaia DR3 211413206732905600 2024.865 0.185 0.130
Gaia DR3 2116748243555689728 2025.4397 0.240 0.160
Gaia DR3 2116748247853607168 2025.4397 0.130 0.090
Gaia DR3 2116892318236994816 2025.4398 0.110 0.080
This is a triple system. Gaia DR3 2119132916774654080 2025.4398 0.180 0.125
@@
This is a triple system. Gaia DR3 2119132916774654208 2025.4398 0.555 0.365
@@
Gaia DR3 2126777988630768000 2025.44 0.375 0.220
Gaia DR3 2126777992924832128 2025.44 0.500 0.325
Gaia DR3 2128304179486008704 2025.4426 0.185 0.130
Gaia DR3 2128304183783972992 2025.4426 0.140 0.100
Gaia DR3 2129540099271049216 2025.4426 0.410 0.230
Gaia DR3 2129540103571086592 2025.4426 0.495 0.280
Gaia DR3 2130903047014230528 2025.4426 0.315 0.185
Gaia DR3 2130903223113004032 2025.4426 0.165 0.115
Gaia DR3 2137630172689997824 2025.4372 0.155 0.105
Gaia DR3 2137630172690002176 2025.4372 0.145 0.100
Gaia DR3 2140767319879292160 2025.4372 0.195 0.140
Gaia DR3 2140767560397457536 2025.4453 0.145 0.105
Gaia DR3 2148617626742496512 2025.4371 0.145 0.100
Gaia DR3 2148617729821714688 2025.4371 0.215 0.145
Gaia DR3 2157622725756418816 2025.4398 0.145 0.105
Gaia DR3 2157623451608162688 2025.4398 0.440 0.235
Gaia DR3 2171870747199482496 2024.6864 0.515 0.295
Gaia DR3 2171870751495498752 2024.6864 0.335 0.190
Gaia DR3 2179261496856372224 2024.6864 0.465 0.290
Gaia DR3 2179261496856372352 2024.6864 0.460 0.280
Gaia DR3 2182397334073261568 2024.6782 0.515 0.310
Gaia DR3 2182397338377483264 2024.6782 0.420 0.235
Gaia DR3 2186439383632375296 2024.6863 1.200 0.885
Gaia DR3 2186439482412215808 2024.6863 0.310 0.200
Gaia DR3 219593745044391552 2024.8618 0.185 0.130
Gaia DR3 219605599154126976 2024.8618 0.245 0.160
Gaia DR3 2196841961921480960 2024.6864 0.430 0.265
Gaia DR3 2196841966222594944 2024.6864 0.875 0.700
Gaia DR3 2201661091331626752 2024.6867 0.125 0.090
Gaia DR3 2201661297490052096 2024.6867 0.120 0.085
Gaia DR3 2201834466266276864 2024.6867 0.115 0.080
Gaia DR3 2201834470572701952 2024.6867 0.140 0.100
Gaia DR3 222299432702445184 2024.8618 0.725 0.535
Gaia DR3 222299436999652096 2024.8618 0.740 0.565
Gaia DR3 2242465891977167872 2025.4426 0.120 0.085
Gaia DR3 2242465891977204992 2025.4426 0.215 0.150
Gaia DR3 229737942400057728 2024.8618 0.295 0.180
Gaia DR3 229737942400695424 2024.8618 0.875 0.675
Gaia DR3 2503402743097074304 2024.8618 0.505 0.290
Gaia DR3 2503402747391938688 2024.8618 0.415 0.225
Gaia DR3 2512615551025883136 2024.6868 0.475 0.270
Gaia DR3 2512615555321101696 2024.6868 0.155 0.110
Gaia DR3 2514529598906714880 2024.8618 0.260 0.165
Too weak Gaia DR3 2514541414361083648 2024.8618 >2.>2. >2.>2.
Gaia DR3 2524864836409161728 2024.6868 0.180 0.130
Gaia DR3 2529474126591876096 2024.6815 0.205 0.140
Gaia DR3 2529661799483080192 2024.6815 0.145 0.105
Gaia DR3 2531509253895359744 2024.6868 0.115 0.085
Gaia DR3 2531509253895360000 2024.6868 0.145 0.105
Gaia DR3 2555905080453468544 2024.6815 0.150 0.105
Gaia DR3 2555905080453468800 2024.6815 0.130 0.095
Gaia DR3 2556421335522520576 2024.6816 0.145 0.105
Gaia DR3 2556421610401138176 2024.6816 0.220 0.150
Gaia DR3 2565584802867037696 2024.6868 0.095 0.065
Gaia DR3 2565584837226776576 2024.6868 0.105 0.070
Gaia DR3 2572437745309855488 2024.6869 0.115 0.070
Gaia DR3 2572437749605558272 2024.6869 0.175 0.120
Gaia DR3 2573278051366910336 2024.6869 0.080 0.060
Gaia DR3 2573278120086386432 2024.6869 0.080 0.060
Triple system, variable Gaia DR3 2581429521337662464 2024.6868 0.130 0.090
Triple system, variable Gaia DR3 2581429521338139264 2024.6868 0.140 0.100
Gaia DR3 2586004584926025600 2024.6869 0.150 0.105
Gaia DR3 2586004589220994560 2024.6869 0.140 0.100
Gaia DR3 2611384547405130752 2024.684 0.170 0.120
Gaia DR3 2611384547405183488 2024.684 0.160 0.115
Gaia DR3 2621082995876646400 2024.684 0.155 0.110
Gaia DR3 2621082995876646528 2024.684 0.185 0.130
Gaia DR3 2638615189818036608 2024.684 0.065 0.055
Gaia DR3 2638615189818036736 2024.684 0.065 0.055
Gaia DR3 2642922251741817344 2024.684 0.255 0.160
Gaia DR3 2642922286101533696 2024.684 0.280 0.170
Gaia DR3 266993755737826176 2024.865 0.535 0.325
Gaia DR3 266993755740665600 2024.865 0.545 0.340
Gaia DR3 2683565561622947584 2024.6839 0.240 0.160
Gaia DR3 2683565630342923520 2024.6839 0.120 0.085
Gaia DR3 2688137537130343936 2024.6838 0.085 0.060
Gaia DR3 2688137537130451456 2024.6838 0.110 0.080
Gaia DR3 2722918766408972288 2024.6839 0.170 0.120
Gaia DR3 2722918796473758720 2024.6839 0.190 0.130
Gaia DR3 2725049173267026560 2024.6838 0.085 0.060
Gaia DR3 2725049173267027968 2024.6838 0.145 0.105
Gaia DR3 2731572610114153856 2024.684 0.095 0.070
Gaia DR3 2731572610114154112 2024.684 0.120 0.090
Gaia DR3 2738818666619511808 2024.6868 0.160 0.110
Gaia DR3 2738818666619512064 2024.6868 0.175 0.120
Gaia DR3 2739496687336986112 2024.6841 0.100 0.070
Gaia DR3 2739498199165473792 2024.6841 0.125 0.090
Gaia DR3 2741770924059034624 2024.6814 0.240 0.150
Gaia DR3 2741770958420224000 2024.6814 0.105 0.075
Gaia DR3 2743364563085132800 2024.6841 0.865 0.710
Gaia DR3 2743364563085132928 2024.6841 0.885 0.720
Gaia DR3 2752042939643757952 2024.6868 0.115 0.080
Gaia DR3 2752042939643758080 2024.6868 0.110 0.075
Gaia DR3 2763789469038658432 2024.6841 0.115 0.080
Gaia DR3 2763789469039961472 2024.6841 0.085 0.060
Too weak Gaia DR3 2799550942998190592 2024.6815 >2.>2. >2.>2.
Gaia DR3 2799550947293531648 2024.6814 0.210 0.140
Gaia DR3 280403536989156224 2024.865 0.160 0.115
Too weak Gaia DR3 280403541287570176 2024.865 >2.>2. >2.>2.
Gaia DR3 2813290096701231104 2024.684 0.385 0.195
Gaia DR3 2813290100998398464 2024.684 0.375 0.200
Gaia DR3 2833895636096202752 2024.684 0.160 0.110
Gaia DR3 2833895636096692224 2024.6757 0.140 0.105
Gaia DR3 2836313908842872704 2024.6866 0.145 0.095
Gaia DR3 2836313908842872832 2024.6866 0.125 0.085
Gaia DR3 2837991381336616320 2024.6867 0.225 0.155
Gaia DR3 2837991454349643136 2024.6867 0.225 0.150
Gaia DR3 2847357575430420992 2024.6815 0.130 0.090
Gaia DR3 2847357575430421120 2024.6815 0.325 0.190
Gaia DR3 2849031405789315328 2024.6814 0.100 0.065
Gaia DR3 2849031410085904640 2024.6814 0.165 0.115
Gaia DR3 2860343426229962752 2024.6814 0.120 0.080
Gaia DR3 2865611564396936192 2024.6866 0.145 0.105
Gaia DR3 2865611564396936320 2024.6866 0.175 0.120
Gaia DR3 2868784063464098048 2024.6866 0.390 0.245
Gaia DR3 2868784136478476032 2024.6866 0.100 0.065
Gaia DR3 2874196929143032192 2024.6815 0.610 0.415
Gaia DR3 2874196929143032448 2024.6815 0.600 0.385
Gaia DR3 2879160021551346176 2024.6866 0.260 0.170
Gaia DR3 2879160021551346688 2024.6866 0.215 0.145
Gaia DR3 293553975230799104 2024.6869 0.155 0.105
Gaia DR3 293554074014451200 2024.6869 0.105 0.065
Gaia DR3 2989629982019538304 2024.8621 1.260 0.925
Gaia DR3 2989630016379276160 2024.8621 1.115 0.930
Gaia DR3 3002343871195786496 2024.8622 0.190 0.135
Gaia DR3 3002531548383047296 2024.8622 0.185 0.130
Gaia DR3 3015129546453825920 2024.8622 0.690 0.475
Gaia DR3 3015130405447132288 2024.8621 0.670 0.445
Gaia DR3 3033070586928188416 2024.8624 0.235 0.160
Gaia DR3 3033070655647665536 2024.8624 0.735 0.580
Gaia DR3 3033419235197542912 2024.8624 0.370 0.215
Gaia DR3 3033442909056369792 2024.8624 0.315 0.195
Gaia DR3 3033693700785580800 2024.8624 0.475 0.280
@@
Gaia DR3 3035042522369773824 2024.8624 0.635 0.435
Gaia DR3 3035042526672692096 2024.8624 0.560 0.415
Gaia DR3 3042299715729341056 2024.8624 0.515 0.315
Gaia DR3 3042300093686461312 2024.8624 0.475 0.285
Gaia DR3 3046204150242468480 2025.2173 0.290 0.185
Gaia DR3 3046204180298475264 2025.2173 0.215 0.150
Gaia DR3 3048356684771809536 2024.8624 0.185 0.135
@@
Gaia DR3 3061488839331542912 2024.8624 0.450 0.260
Gaia DR3 3061488839334653696 2024.8624 0.455 0.270
Gaia DR3 3075884367113894400 2024.8625 0.785 0.650
Gaia DR3 3075884367113894528 2024.8625 0.610 0.415
Gaia DR3 3079739250589041024 2024.8625 0.850 0.650
Gaia DR3 3079739254881696256 2024.8625 0.590 0.340
Gaia DR3 3084652044636398976 2024.8625 0.175 0.125
Gaia DR3 3084652044636399104 2024.8625 0.190 0.135
Gaia DR3 3103862814877943296 2024.8622 0.190 0.135
Gaia DR3 3103862814877944064 2025.2174 0.225 0.155
Gaia DR3 3107726150841412224 2024.8624 0.770 0.655
Gaia DR3 3107726155141310080 2024.8624 0.760 0.625
Gaia DR3 3155908468358227968 2024.8624 0.600 0.425
Gaia DR3 3155908575735398400 2024.8624 0.505 0.335
Gaia DR3 3158878734598549376 2025.2173 0.240 0.155
Gaia DR3 3158878734598549504 2024.8623 0.315 0.185
Gaia DR3 3158926322836178816 2025.2173 0.210 0.140
Gaia DR3 3158926322836178944 2024.8623 0.475 0.245
Gaia DR3 3170300942420466176 2025.2173 0.140 0.100
Gaia DR3 3170303755624614528 2024.8623 0.240 0.155
Gaia DR3 3170303759920420224 2024.8623 0.165 0.115
Gaia DR3 3170394607068638336 2025.2173 0.150 0.110
Gaia DR3 3174712286151938944 2024.8621 0.595 0.400
Gaia DR3 3174712354871414912 2024.8621 0.485 0.290
Gaia DR3 3194331181363949312 2024.8619 0.480 0.275
Gaia DR3 3194331284443164288 2024.8619 0.430 0.240
@@
Gaia DR3 3194720889516362880 2024.8619 0.540 0.360
Gaia DR3 3201928153876092416 2024.862 0.890 0.730
Gaia DR3 3201928153876683648 2024.862 0.985 0.800
Gaia DR3 3208097204381858304 2024.8622 0.390 0.215
Gaia DR3 3208097582339739136 2024.8622 0.355 0.200
Gaia DR3 3208920321978859008 2024.8622 0.215 0.150
Gaia DR3 3208920326274769664 2024.8622 0.740 0.535
@@
Gaia DR3 3211638593896898944 2024.8622 0.210 0.140
Gaia DR3 3213811602532945152 2024.8622 0.740 0.565
Gaia DR3 3213811641189319680 2024.8622 0.760 0.570
Gaia DR3 3216741186843662720 2024.8622 1.000 0.800
Gaia DR3 3216741186843662848 2024.8622 1.440 1.220
@@
Gaia DR3 3223175082210856192 2024.8622 0.685 0.440
Gaia DR3 3233274852427241216 2024.862 0.255 0.165
Too weak Gaia DR3 3233274955506455936 2024.862 >2.>2. >2.>2.
Gaia DR3 3237157051889659904 2024.8622 0.350 0.200
Gaia DR3 3237157159265190400 2024.8622 0.415 0.230
Gaia DR3 3239798461074336512 2024.8649 0.340 0.195
Gaia DR3 3239798461074336640 2024.8649 0.395 0.215
Gaia DR3 3242368702645853184 2024.8649 0.415 0.225
Gaia DR3 3242368706941778688 2024.8649 0.385 0.215
Gaia DR3 3251884739561945472 2024.862 1.450 1.170
Gaia DR3 3251884739561945600 2024.862 0.465 0.270
Gaia DR3 3258912371210481152 2024.862 0.360 0.205
Gaia DR3 3258912439929957376 2024.862 0.595 0.415
Gaia DR3 3266980170921153920 2024.8618 0.590 0.370